# Physics-Informed Recurrent Network for Gas Pipeline Network Parameters Identification

### v1

@@ 3.4 Training Techniques of Physics-Informed Recurrent Network
(2) Data normalization
The substantial disparities in pressure and mass flow rate within the gas network, e.g., $10^{5}$-$10^{6}$ and $10^{0}$-$10^{1}$, pose challenges for training physics-informed recurrent network model. By employing data normalization techniques, diverse physical quantities can be normalized to comparable magnitudes, ensuring numerical stability throughout the training process.
(3) Gradient clipping and truncated backward propagation through time method
Similar to the training process of the standard RNN, the backward propagation through time process also involves a chain of matrix products. High matrix powers may lead to numerical instability, causing gradients to either explode or vanish [36]. The gradient clipping method can be used to solve this problem. In each time step, the gradient is checked and forced to the upper bound or lower bound if it exceeds the limits. Besides, the truncated backward propagation through time method [35] can be applied. This method segments long-term consequences into several shorter ones, reducing the highest matrix powers, thereby enhancing model simplicity and stability.
4 Numerical Tests
4.1 Simulation Setup
The case studies are carried out on four different scales of gas network topologies from the GasLib dataset [37], including GasLib-11, GasLib-24, GasLib-40, and GasLib-135 topologies.
The programs are implemented on the Python 3.11 platform using the PyTorch framework [38]. The RMSprop optimizer is employed, with a learning rate ranging from $10^{-5}$ to $10^{-3}$, depending on the specific test cases and quality of training data. A learning rate scheduler is applied with a step size of 4, while the decay factor (gamma) is set to 0.2. The batch size is configured to 16. All models are trained on a Tesla P100 GPU with 16 GB of VRAM.
@@ 4.1 Simulation Setup
4.2 Performance of Pipelines Parameters Identification
To evaluate the performance of the proposed PIRN model, the GasLib-24 system is used as a test case for identifying pipeline parameters. During the warm start process, the pipeline parameters requiring estimation are initialized with values that contain estimation errors. In this numerical test, three levels of estimation error—within the ranges of ±5%, ±15%, and ±25%—are introduced to assess the model’s performance. Random errors are added to the initial estimates within these ranges. The model is then trained for 20 epochs, using measurement data from 100 batches. At the start of each epoch, the data sequence is reshuffled to enhance learning. Figure 6 displays the identification process on a logarithmic scale for different initial estimation errors.
Figure 6: Parameter identification process of PIRN with different initial estimation errors.
Based on the results presented in Figure 6, the parameter identification process successfully converges across different ranges of initial estimation errors. To assess the performance of the identified state-space model for the gas pipeline network, we employed our model to compute the gas states at each node. These calculated results were then compared to the FDM simulation results. The accuracy was quantified using the mean average percentage error (MAPE) indicator, as presented in Table 2. The results demonstrate that the PIRN model effectively identifies the pipeline parameters, despite varying levels of initial estimation errors.
Table 2: Parameters Identification Results
@@ 4.2 Performance of Pipelines Parameters Identification
4.3 Robustness under Measurement Data with Outliers
In the PIRN model, we employ regularization, gradient clipping, mini-batch stochastic gradient descent, and learning rate scheduling techniques to mitigate the effect of measurement outliers during training. To assess its robustness, outliers with an amplitude of 15% at varying proportions are randomly introduced to simulate different noise levels in gas network data. Our parameter identification method is then applied to analyze the noise-affected data and identify gas network parameters. Figure 7 illustrates the parameter identification process of PIRN under different noise levels in the GasLib-24 system. As measurement noise deviations increase, the parameters can temporarily be trapped in local optima due to incorrect gradient descent directions affected by measurement outliers, as shown by the 5% noise curve in Figure 7. Nevertheless, the model ultimately converges after additional training iterations.
Figure 7: Parameter identification process of PIRN under different proportions of measurement outliers.
In addition, the MAPE of parameters identification results under different measurement noise with different gas pipeline network systems are listed in Table 3.
Table 3: Parameters Identification Results under Different Proportions of Outliers
@@ 4.3 Robustness under Measurement Data with Outliers
4.4 Parameter Efficiency Compared to Standard Neural Networks
The proposed PIRN integrates the dynamic state-space function of the gas pipeline network, resulting in a model with an explicit matrix operation form and a relatively small number of required parameters to achieve strong fitting performance. In this test case, we compare the PIRN model with the classic Elman RNN and LSTM networks. Both of these networks are trained using control variables $\bm{u}_{t}$ as inputs and measurement data $\bm{y}_{t}$ as output labels.
@@ 4.4 Parameter Efficiency Compared to Standard Neural Networks
According to Figure 8, under the condition of achieving similar fitting accuracy, both Elman RNN and LSTM networks require significantly more parameters compared to the proposed PIRN. Moreover, as the network scale increases, this disparity becomes increasingly pronounced. Therefore, compared to RNN and LSTM networks, the proposed PIRN demonstrates better parameter efficiency compared to the standard neural networks.
4.5 Integrate the Identified State-Space Model in the Gas-Electricity Integrated Optimal Dispatch Problem
The PIRN-based parameter identification method produces an explicit state-space function for the gas pipeline network, which can directly integrate in the optimization problem. In this subsection, the IEEE RTS96 One Area 24-bus power system and the 24-pipe benchmark gas system [39] are used as numerical test cases to demonstrate how the identification results can be integrated into the optimal dispatch problem. The gas system includes 4 GTs, 4 P2Gs, and 5 compressors, while the power system consists of 1 coal generator, 3 wind turbines, and 2 photovoltaic units. Since modeling of gas-electricity integrated optimal dispatch is not the main focus of this paper, the detail descriptions of the system configurations and mathematical models can be found in the A and B, respectively. The total power system load and renewable generation curves are presented in Figure 9.
@@ 4.5 Integrate the Identified State-Space Model in the Gas-Electricity Integrated Optimal Dispatch Problem
Figure 10 and Figure 11 displays the gas generation curves of P2Gs and gas consumption of GTs throughout the day, respectively. During three specific timeframes (2:00-3:00, 7:00-10:00, and 11:30-13:00), the output power from renewable energy generators surpasses the total power load of the system. Consequently, P2Gs undergo electrolysis to produce natural gas for energy storage purposes. During these intervals, GTs stop gas consumption for electricity generation.
@@
Figure 12: Gas pressures of serval terminal nodes throughout the day.
Figure 13: Gas flow rates of serval pipelines throughout the day.
Figure 12 and Figure 13 displays the gas mass flow rates in multiple pipes and the gas pressures at various nodes throughout the day, respectively. The solid lines represent our model’s calculated results, contrasting with the dashed lines reflecting outcomes from the FDM results. Then, we use the statistical indicators MAPE, RMSE and $R^{2}$ to demonstrate the performance of our method. Error statistics of gas pressure and gas mass power flow throughout the day are shown in Table 4. The identified state-space model can represent the dynamic process of the gas system with high calculation accuracy. Therefore, it can be effectively embedded in the scenarios such as optimal dispatch of integrated energy systems.
Table 4: Error Statistics of Gas Pressure and Gas Mass power Flow
@@ 5 Conclusion
This paper introduces a physics-informed recurrent network model to identify the state-space model of the gas pipeline network using sparse measurement data from terminal nodes of pipelines. It combines data-driven learning from measurements with fluid partial differential equations governing the dynamic states of the gas pipeline network. This physics-informed recurrent network incorporates a physical embedding by substituting artificial neural perceptron units with the gas pipeline network’s physical state-space model, ensuring physical interpretability.
@@
Utilizing the PyTorch framework, the parameters of the state-space model can be accurately learned. Case studies demonstrate that our proposed framework accurately identifies the gas pipeline network model and simultaneously estimates the internal states of gas along the pipelines. Additionally, the identified state-space model can be conveniently applied in the joint optimal dispatch of the integrated gas-electricity system problem and it performs high accuracy and parameter efficiency.
@@
The PIRN framework uses far fewer parameters than standard neural networks, making it more sensitive to individual parameter values. This heightened sensitivity presents a challenge in quickly determining optimal hyper-parameters to improve convergence speed, requiring further exploration. Additionally, extending the PIRN framework to more complex models of gas pipeline networks, such as nonlinear models, is another important avenue for future research.
Acknowledgments
This work was supported by the Science-Technique Plan Project of Jiangsu Province (No. SBE2022110130) and the Beijing Natural Science Foundation (L243003).
Appendix A Parameters of Numerical Tests
In our numerical tests, the IEEE RTS96 One Area 24 bus power system and 24 pipe benchmark gas system [39] are used as the gas-electricity integrated optimal dispatch test case. The parameters of GTs, P2Gs and compressors in the gas network and generators, wind turbines and photovoltaics in the power system are listed in the Table 5, 6 and 7.
Table 5: Parameters of GTs and P2Gs
Type of Units Bus in power system Node in gas system Total capacity (MW) Energy efficiency
GTs 6 8 60 0.75
GTs 13 12 80 0.77
GTs 19 18 70 0.72
GTs 25 24 60 0.69
P2Gs 4 6 8 0.48
P2Gs 9 13 7 0.44
P2Gs 3 19 9 0.50
P2Gs 22 25 6 0.54
Table 6: Parameters of Compressors
Compressors Pipeline in gas system Maximum additional gas pressure (MPa) Locating distance from the head node (km)
COMP1 P1 0.1 2
COMP2 P8 0.05 5
COMP3 P3 0.03 1
COMP4 P14 0.04 3
COMP5 P20 0.06 1
Table 7: Parameters of Generators in the Power System
Type of unites Bus in power system Total capacity (MW)
Coal generators 2 100
Wind turbines 4 65
Wind turbines 18 80
Wind turbines 21 55
Photovoltaics 16 50
Photovoltaics 22 70
Appendix B Mathematical Model of Gas-electricity Integrated Optimal Dispatch
B.1 Models of Gas System
According to the parameter identification results, the state-space model of the gas pipeline network is as follows:
$\bm{h}_{t}=\bm{J}^{(1)}\bm{S}\bm{h}_{t-1}+\bm{J}^{(2)}\bm{u}_{t}$ (36)
where the matrices $\bm{J}^{(1)}$ and $\bm{J}^{(2)}$ can be obtained by the parameter identification model. $\bm{u}_{t}$ denotes the control variable vector, collecting all the controllable gas pressure and gas mass flow injection of nodes.
The gas wells (GWs) provide natural gas for the whole gas system with the upper and lower gas injection and gas pressure bounds.
$\underline{f}_{k}^{\mathrm{gw}}\leq f_{k,t}^{\mathrm{gw}}\leq\overline{f}_{k}^{\mathrm{gw}}$ (37)
$\underline{\pi}_{k}^{\mathrm{gw}}\leq\pi_{k,t}^{\mathrm{gw}}\leq\overline{\pi}_{k}^{\mathrm{gw}}$ (38)
where $\pi_{k,t}^{\mathrm{gw}}$ and $f_{k,t}^{\mathrm{gw}}$ denote the gas pressure and gas mass flow of gas well $k$ at time $t$, respectively.
The compressors also have the maximum and minimum pressure booster limits.
$\Delta\underline{\pi}_{k}\leq\Delta\pi_{k,t}\leq\Delta\overline{\pi}_{k}$ (39)
where $\Delta\pi_{k,t}$ denotes the boosted gas pressure of compressor $k$ at time $t$.
The gas turbines (GTs) transform the natural gas into electricity with the efficiency $\eta_{k}^{\mathrm{gt}}$. Both their gas input and the electricity output have the upper and lower bounds.
$p_{k,t}^{\mathrm{gt}}=\eta_{k}^{\mathrm{gt}}f_{k,t}^{\mathrm{gt}}q_{L}$ (40)
$\underline{f}_{k}^{\mathrm{gt}}\leq f_{k,t}^{\mathrm{gt}}\leq\overline{f}_{k}^{\mathrm{gt}}$ (41)
where $p_{k,t}^{\mathrm{gt}}$ and $f_{k,t}^{\mathrm{gt}}$ denote the generation power and gas mass flow of gas turbine $k$ at time $t$, respectively. $\eta_{k}^{\mathrm{gt}}$ denotes the efficiency of gas-power transformation. $q_{L}$ denotes the low heat value of natural gas.
The power-to-gas devices (P2Gs) transform the electricity into natural gas with the efficiency $\eta_{k}^{\mathrm{p2g}}$. Both their electricity input and gas output have the upper and lower bounds.
$f_{k,t}^{\mathrm{p2g}}=\eta_{k}^{\mathrm{p2g}}p_{k,t}^{\mathrm{p2g}}/q_{L}$ (42)
$\underline{f}_{k}^{\mathrm{p2g}}\leq f_{k,t}^{\mathrm{p2g}}\leq\overline{f}_{k}^{\mathrm{p2g}}$ (43)
where $f_{k,t}^{\mathrm{p2g}}$ and $p_{k,t}^{\mathrm{p2g}}$ denote the generated gas mass flow and power consumption of P2G $k$ at time $t$, respectively.
For the consideration of security, the gas pressures of gas pipelines also have upper and lower bounds. For any $l\in\mathcal{L}$
$\underline{\pi}_{l}\leq\pi_{i,t}\leq\overline{\pi}_{l},\forall i\in\mathcal{S}_{l}$ (44)
where $\pi_{i,t}$ denotes the gas pressure of gas pipeline at the discrete computation point $i$ at time $t$. $\mathcal{L}$ denotes the index set of pipelines. $\mathcal{S}_{l}$ denotes the index set of discrete computation point of pipeline $l$.
The gas injection at nodes can be described as the net injection of all the units. For any $n\in\mathcal{N}$
$f_{n,t}^{\mathrm{inj}}=\sum_{k\in\mathcal{D}_{n}^{\mathrm{gw}}}f_{k,t}^{\mathrm{gw}}+\sum_{k\in\mathcal{D}_{n}^{\mathrm{p2g}}}f_{k,t}^{\mathrm{p2g}}-\sum_{k\in\mathcal{D}_{n}^{\mathrm{gt}}}f_{k,t}^{\mathrm{gt}}-f_{n,t}^{\mathrm{ld}}$ (45)
where $\mathcal{D}_{n}^{\mathrm{gw}}$, $\mathcal{D}_{n}^{\mathrm{p2g}}$, and $\mathcal{D}_{n}^{\mathrm{gt}}$ denote the index sets of gas sources, P2Gs, and GTs that directly linked to node $n$, respectively. $f_{n,t}^{\mathrm{ld}}$ denotes the mass flow rate of gas load at node $n$ at time $t$. $\mathcal{N}$ denotes the index set of all pipeline nodes.
B.2 Models of Power System
In the power system, there are some units such as GTs, P2Gs, generators (GENs), wind turbines (WTs) and photovoltaics (PVs). Each type of unit has its specific active and reactive power flexibility range and can be express as a polygon. For any $s=\{\mathrm{gt,p2g,gen,wt,pv}\}$ and any $k\in\mathcal{K}^{s}$
$\underline{p}_{k}^{s}\leq p_{k,t}^{s}\leq\overline{p}_{k}^{s}$ (46)
$\underline{q}_{k}^{s}\leq q_{k,t}^{s}\leq\overline{q}_{k}^{s}$ (47)
where $\mathcal{K}^{s}$ denotes the index sets of unit $s$ in the power system. $p_{k,t}^{s}$ and $q_{k,t}^{s}$ denote the active and reactive output power of unit $k$-th at time $t$.
The ramp rate constraints should of coal generators should also be considered as follows:
$-r_{k}^{\mathrm{gen}}\leq p_{k,t}^{\mathrm{gen}}-p_{k,t-1}^{\mathrm{gen}}\leq r_{k}^{\mathrm{gen}}$ (48)
where $r_{k}^{\mathrm{gen}}$ denotes the ramp rates.
The network constraints of the power system restrict the active and reactive output power among all the units. Based on the linearized power flow model, the network constraints can be expressed in the compact form as follows:
$\bm{A}_{t}^{\mathrm{ps}}\bm{x}_{t}\leq\bm{b}_{t}^{\mathrm{ps}}$ (49)
where $\bm{A}_{t}^{\mathrm{ps}}$ and $\bm{b}_{t}^{\mathrm{ps}}$ are constant parameters representing the network constraints of the power system; $\bm{x}_{t}$ is a vector collects the active and reactive output power of all units at time $t$, that is
[…]
B.3 Objective Function
The objective of optimal dispatch of integrated gas-electricity system is minimizing the total cost of these two systems.
$\min\sum_{t\in\mathcal{T}}\left(\sum_{i\in\mathcal{K}^{\mathrm{gen}}}C_{i}^{\mathrm{gen}}+\sum_{j\in\mathcal{D}^{\mathrm{gt}}}C_{j}^{\mathrm{gt}}\right)\cdot\Delta t$ (51)
$C_{i}^{\mathrm{gen}}=a_{i}^{\mathrm{gen}}\left(p_{i,t}^{\mathrm{gen}}\right)^{2}+b_{i}^{\mathrm{gen}}p_{i,t}^{\mathrm{gen}}+c_{i}^{\mathrm{gen}}$ (52)
$C_{j}^{\mathrm{gt}}=\lambda_{j}^{\mathrm{gt}}f_{j,t}^{\mathrm{gt}}$ (53)
where $C_{i}^{\mathrm{gen}}$ and $C_{j}^{\mathrm{gt}}$ denote the fuel cost of $i$-th generator and $j$-th GT, respectively; $a_{i}^{\mathrm{gen}}$, $b_{i}^{\mathrm{gen}}$ and $c_{i}^{\mathrm{gen}}$ are coefficients of cost function of $i$-th generator. $\lambda_{j}$ are cost of natural gas of $j$-th GT. Since the renewable generators such as wind turbines and photovoltaics do not need to consume fuel, their costs are set to 0.

### v2

@@ III-D1 Warm Start
III-D2 Data Normalization
The substantial disparities in pressure and mass flow rate within the gas network, e.g., $10^{5}$-$10^{6}$ and $10^{0}$-$10^{1}$, pose challenges for training PIRN model. By employing data normalization techniques, diverse physical quantities can be normalized to comparable magnitudes, ensuring numerical stability throughout the training process.
III-D3 Gradient Clipping and Truncated Backward Propagation Through Time Method
Similar to the training process of the classic RNN, the backward propagation through time process also involves a chain of matrix products. High matrix powers may lead to numerical instability, causing gradients to either explode or vanish. The gradient clipping method can be used to solve this problem. In each time step, the gradient is checked and forced to the upper bound or lower bound if it exceeds the limits. Besides, the truncated backward propagation through time method [26] can be applied. This method segments long-term consequences into several shorter ones, reducing the highest matrix powers, thereby enhancing model simplicity and stability.
IV Numerical Tests
IV-A Simulation Setup
The case studies are carried out on four gas-network topologies of varying scale from the GasLib dataset [27], including GasLib-11, GasLib-24, GasLib-40, and GasLib-135 topologies.
The programs are implemented on the Python 3.11 platform using the PyTorch framework. The RMSprop optimizer is employed, with a learning rate ranging from $10^{-5}$ to $10^{-3}$, depending on the specific test cases and quality of training data. A learning rate scheduler is applied with a step size of 4, while the decay factor (gamma) is set to 0.2. The batch size is configured to 16. All models are trained on a Tesla P100 GPU with 16 GB of VRAM.
@@ IV-A Simulation Setup
IV-B Performance of Pipelines Parameters Identification
To evaluate the performance of the proposed PIRN model, the GasLib-24 system is used as a test case for identifying pipeline parameters. During the warm-start process, the pipeline parameters requiring estimation are initialized with values that contain estimation errors. In this numerical test, three levels of estimation error, within the ranges of ±5%, ±15%, and ±25%, are introduced to assess the model’s performance. Random errors are added to the initial estimates within these ranges. The model is then trained for 20 epochs, using measurement data from 100 batches. At the start of each epoch, the data sequence is reshuffled to enhance learning. Fig. 6 displays the identification process on a logarithmic scale for different initial estimation errors.
Figure 6: Parameter identification training process of PIRN with different initial estimation errors.
Based on the results presented in Fig. 6, the parameter identification process successfully converges across different ranges of initial estimation errors. To assess the performance of the identified state-space model for the gas pipeline network, we employed our model to compute the gas states at each node. These calculated results were then compared to the FDM simulation results. The accuracy was quantified using the mean average percentage error (MAPE) indicator, as presented in TABLE II. The results demonstrate that the PIRN model effectively identifies the pipeline parameters, despite varying levels of initial estimation errors.
TABLE II: Parameters Identification Results
@@ IV-B Performance of Pipelines Parameters Identification
IV-C Robustness under Measurement Data with Outliers
In the PIRN model, we employ regularization, gradient clipping, mini-batch stochastic gradient descent, and learning rate scheduling techniques to mitigate the effect of measurement outliers during training. To assess its robustness, outliers with an amplitude of 15% at varying proportions are randomly introduced to simulate different noise levels in gas network data. Our parameter identification method is then applied to analyze the noise-affected data and identify gas network parameters. Fig. 7 illustrates the parameter identification process of PIRN under different noise levels in the GasLib-24 system. As measurement noise deviations increase, the parameters can temporarily be trapped in local optima due to incorrect gradient descent directions affected by measurement outliers, as shown by the 5% noise curve in Fig. 7. Nevertheless, the model ultimately converges after additional training iterations.
Figure 7: Parameter identification training process of PIRN under different proportions of measurement outliers.
In addition, the MAPE of parameters identification results under different measurement noise with different gas pipeline network systems are listed in TABLE III.
TABLE III: Parameters Identification Results under Different Proportions of Outliers
@@ IV-C Robustness under Measurement Data with Outliers
IV-D Parameter Efficiency Compared to Classic Neural Networks
The proposed PIRN integrates the dynamic state-space function of the gas pipeline network, resulting in a model with an explicit matrix operation form and a relatively small number of required parameters to achieve strong fitting performance. In this test case, we compare the PIRN model with the classic Elman RNN and LSTM networks. All networks are trained using control variables $\bm{u}_{t}$ as inputs and measurement data $\bm{y}_{t}$ as output labels.
@@ IV-D Parameter Efficiency Compared to Classic Neural Networks
According to Fig. 8, under the condition of achieving similar fitting accuracy, both Elman RNN and LSTM networks require significantly more parameters compared to the proposed PIRN. Moreover, as the network scale increases, this disparity becomes increasingly pronounced. Therefore, compared to RNN and LSTM networks, the proposed PIRN demonstrates better parameter efficiency compared to the classic neural networks.
IV-E Integrate the Identified State-Space Model in the Gas-Electricity Integrated Optimal Dispatch Problem
The PIRN-based parameter identification method produces an explicit state-space function for the gas pipeline network as in (26), which can directly integrate in the gas-electricity integrated optimal dispatch problem. In this subsection, the IEEE RTS96 One Area 24-bus power system and the 24-pipe benchmark gas system are used as numerical test cases to demonstrate how the identification results can be integrated into the optimal dispatch problem. The detailed configuration and topology of this system can be found in [28]. The gas system includes 4 GTs, 4 P2Gs, and 5 compressors, while the power system consists of 1 coal generator, 3 wind turbines, and 2 photovoltaic units. Since modeling of gas-electricity integrated optimal dispatch is not the main focus of this paper, the detailed descriptions of the system configurations and mathematical models can be found in the supplementary file. The total power system load and renewable generation curves are presented in Fig. 9.
@@ IV-E Integrate the Identified State-Space Model in the Gas-Electricity Integrated Optimal Dispatch Problem
Fig. 10 and Fig. 11 display the gas generation curves of P2Gs and gas consumption of GTs throughout the day, respectively. During three specific timeframes (2:00-3:00, 7:00-10:00, and 11:30-13:00), the output power from renewable energy generators surpasses the total power load of the system. Consequently, P2Gs undergo electrolysis to produce natural gas for energy storage purposes. During these intervals, GTs stop gas consumption for electricity generation.
@@
Figure 12: Gas pressures of several terminal nodes throughout the day.
Figure 13: Gas flow rates of several pipelines throughout the day.
Fig. 12 and Fig. 13 displays the gas mass flow rates in multiple pipes and the gas pressures at various nodes throughout the day, respectively. The solid lines represent our model’s calculated results, contrasting with the dashed lines reflecting outcomes from the FDM results. Then, we use the statistical indicators MAPE, RMSE and $R^{2}$ to demonstrate the performance of our method. Error statistics of gas pressure and gas mass power flow throughout the day are shown in TABLE IV. The identified state-space model can represent the dynamic process of the gas system with high calculation accuracy. Therefore, it can be effectively embedded in the scenarios such as optimal dispatch of integrated energy systems.
TABLE IV: Error Statistics of Gas Pressure and Gas Mass Power Flow
@@ V Conclusion
This paper introduces a physics-informed recurrent network model to identify the state-space model of the gas pipeline network using sparse measurement data from terminal nodes of pipelines. It combines data-driven learning from measurements with fluid partial differential equations governing the dynamic states of the gas pipeline network. This PIRN incorporates a physical embedding by substituting artificial neural perceptron units with the gas pipeline network’s physical state-space model, ensuring physical interpretability.
@@
Utilizing the PyTorch framework, the parameters of the state-space model can be accurately learned. Case studies demonstrate that our proposed framework accurately identifies the gas pipeline network model and simultaneously estimates the internal states of gas along the pipelines. Additionally, the identified state-space model can be conveniently applied in the joint optimal dispatch of the integrated gas-electricity system problem and it achieves high accuracy and parameter efficiency.
@@
The PIRN framework uses far fewer parameters than classic neural networks, making it more sensitive to individual parameter values. This heightened sensitivity presents a challenge in quickly determining optimal hyper-parameters to improve convergence speed, requiring further exploration. Additionally, extending the PIRN framework to more complex models of gas pipeline networks, such as nonlinear models, is another important avenue for future research.